OpenStudy (asylum15):

Solve the following trigonometric equation for 0 ≤ A ≤ 2π : 1+cos A=sin^2A.

OpenStudy (asylum15):

\[\sin ^{2}A\]

OpenStudy (anonymous):

the standard formula says \[\sin ^{2} A+\cos ^{2} A=1\] so sin ^{2} A=1-cos ^{2} A substituting this in the equation u get 1+cos A= 1-cos ^{2} A => cos ^{2} A+cos A=0 => cos A ( cos A+1)=0 => either cos A = 0 => A = 0 as cos 0 = 0 Or cos A+1=0=>cos A = -1 => A = pi degrees as Cos pi = -1 (it falls in second quadrant where all trigonometric functions are negative other than sine) so A is either 0 or pi

OpenStudy (asylum15):

thanks :D